# Foundation Models Meet Low-Cost Sensors: Test-Time Adaptation for Rescaling Disparity for Zero-Shot Metric Depth Estimation

### v1

@@ IV-C Results
Ours rescaling w/ SFM (SIFT [41]) 0.794 0.188 6.981
Ours rescaling w/ SFM (OmniGlue [42]) 0.947 0.112 5.300
TABLE V: Comparison of the metric depth estimated by SGM [39] or by rescaling Depth Anything predictions using SGM.
Methods KITTI DIML Outdoor
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
SGM [39] 0.855 0.16 7.351 0.973 0.118 1.503
Ours rescaling w/ stereo (SGM [39]) 0.941 0.072 3.526 0.96 0.123 1.509
TABLE VI: Cost comparison for different architectures.
@@
Ours Depth Anything [2] + rescaling 335 120
In Tab. I, we compare our LiDAR-based rescaling approach with zero-shot metric depth estimation methods and other depth estimation methods that have been fine-tuned on the NYU V2 dataset [33]. We conduct a similar study in Tab. II on the KITTI dataset [37] with in addition results of rescaling with 3D reference points provided by structure-from-motion or by stereo depth estimation. For a fair comparison, we also include the performance of stereo depth estimation using Semi Global Matching [39] only. Tab. III and Tab. IV provide additional comparisons on indoor and outdoor datasets, respectively. For each dataset, our results show rescaling using LiDARs with different numbers of beams and for the DDAD dataset [23], we also leverage structure-from-motion. Finally, we analyze the cost of some of the methods we compare in Tab. VI. The runtimes have been measured on the same NVIDIA GeForce RTX $3090$ GPU for each approach.
Rescaling with LiDAR
Our experiments highlight an overall benefit of using our rescaling approach with $32$-laser or $16$-laser LiDAR rather than using zero-shot metric depth estimation. Thus, we observe on average $46$% improvement on $\delta_{1}$, $0.5$% on AbsRel and $47$% on RMSE relative to the other zero-shot methods. In contrast, the methods that have been fine-tuned on the same domain as the test set (see lines with (FT) in Tab. I and Tab. II) compare favorably to ours. These approaches benefit from an additional in-domain training which is often costly to perform due to dataset creation and training computation. We notice that our method is sensitive to the domain with on average $28$% and $13$% enhancements on indoor (including NYU V2) and outdoor (including KITTI) datasets, respectively, relative to other zero-shot methods whatever the metric. Regarding performance with 2D LiDAR, the same sensitivity to the domain is equally apparent. However, the results are a bit lower than zero-shot and fine-tuned metric depth estimation methods and lower than our other approaches, especially on outdoor datasets while remaining competitive on indoor datasets. Interestingly, we observe that increasing the number of beams from 16 to 32 does not necessarily improve the performance.
Rescaling with stereo
We show in our experiments on KITTI dataset that using stereo depth estimation with SGM [39] in order to rescale Depth Anything affine-invariant disparity predictions is more efficient than directly using SGM estimates. This improvement can be explained by the limited range of stereo methods that is exceeded by the depth of field in KITTI images. We carry out the same experiments on the DIML outdoor dataset [38] where the depth of field is much lower than in KITTI. Indeed, the results in Tab. V highlight that directly estimating depth with SGM [39] is better than using these estimates as a reference for rescaling even if DIML is part of the Depth Anything training set. Relative to the other methods we compare on the KITTI benchmark Tab. II, rescaling with stereo does not work as well as fine-tuned methods or rescaling with LiDAR 16 and 32 beams. However, rescaling with stereo performs on par with zero-shot metric depth estimation and is better than rescaling with SFM and LiDAR 1 beam.
Rescaling with structure-from-motion
Our experiments highlight that structure-from-motion techniques can provide reliable reference points for rescaling Depth Anything affine-invariant disparity predictions. More specifically, we show that results with SIFT [41] are slightly lower than the ones of other zero-shot monocular metric depth estimation methods, while using OmniGlue [42] increases performance relative to other zero-shot methods of the $\delta_{1}$, AbsRel and RMSE metrics on average by $7$%, $4$% and $19$%, respectively. Compared to the other approaches, rescaling with SFM performs better than rescaling with LiDAR 1 beam but worse than other methods that may however be more costly. This includes rescaling with LiDAR 16 and 32 beams or fine-tuned approaches that requires an extra training on a dataset that, in real cases, must be created.
Inference cost study
@@
To complete our study, we analyze the inference cost of rescaling Depth Anything V1 affine-invariant disparity maps with other fine-tuned or zero-shot metric depth estimation methods in Tab. VI. For this purpose, we compare two informative and easy-to-access variables which are the number of parameters and the average inference runtime of these approaches for a single image. We notice that zero-shot methods are more costly at inference than our rescaling approach because they introduce extra modules such as a ConvGRU block in [31] or a camera module in [32] that returns a dense representation of the camera calibration. In contrast, fine-tuned methods have a reduced inference cost relative to ours thanks to a more complex training. Finally, as each image is processed independently, the computation overhead of our method can be even reduced when applied to a video stream. Indeed, since the neural network runs on the GPU and the RANSAC on the CPU, the inference of the neural network at step $t+1$ can overlap the execution of RANSAC at step $t$.
IV-D Comparison with Depth Anything V2
@@
TABLE VII: Comparative study between Depth Anything V1 and Depth Anything V2 on the KITTI dataset [37].
Rescaling with Depth Anything KITTI
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$ $R^{2}$uparrow
@@
We compare performance between Depth Anything V1 [2] and Depth Anything V2 [4] on the KITTI dataset in Tab. VII. We notice that Depth Anything V1 [2] slightly outperforms Depth Anything V2 [4] when used in our rescaling approach. Moreover, Depth Anything V1 has a higher coefficient of determination $R^{2}$ that corresponds to the proportion of the variance of the metric disparity that is explainable by the linear regression model parameterized with the $\alpha$ and $\beta$ that has been found. This means that while Depth Anything V2 manages to handle small details better than Depth Anything V1 (see Fig. 2), the latter predicts more accurate disparity maps within an affine transformation.
@@
IV-E Ablation study
TABLE VIII: Ablation study to validate the need for estimating the rescaling parameters for each image
Rescaling with NYU V2
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
@@ IV-E Ablation study
KITTI
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
@@
TABLE IX: Ablation study to validate the use of the RANSAC algorithm.
Methods RANSAC KITTI
$\delta_{1}$uparrow AbsRel $\downarrow$ RMSE $\downarrow$
Ours rescaling w/ lidar 1 beam 0.897 0.113 4.871
@@
Ours rescaling w/ lidar 16 beams 0.966 0.063 2.849
@@
Ours rescaling w/ lidar 32 beams 0.966 0.063 2.852
@@
Ours rescaling w/ stereo (SGM) 0.904 0.086 4.336
0.941 0.072 3.526
Tab. VIII compares dynamic rescaling i.e. an estimation of the parameters $\alpha$ and $\beta$ for each image to a static rescaling with unique parameters for all the images of a dataset. In the latter case, the parameters are the mean $\alpha$ and $\beta$ obtained with the best rescaling (here with LiDAR 16 beams). The results show that performing a rescaling for each image tends to provide better performance even if the camera does not change or if the image domain remains similar.
The ablation study in Tab. IX aims to justify the use of a RANSAC algorithm [6] when estimating the parameters of the affine transformation to recover the metric depth. We observe that the RANSAC is beneficial most of the time, especially when the 3D reference points are not provided by a LiDAR. Indeed, structure-from-motion or stereo-depth techniques are more likely to generate outliers that would disturb a vanilla linear regression but would be filtered out by the RANSAC.
@@ V Conclusion
In this paper, we provide a low-cost method to estimate metric depth based on Depth Anything V1 [2] affine-invariant disparity predictions that are rescaled using 3D reference points provided by an additional sensor. Thus, the solution we propose is adaptable to any camera calibration and cheaper than traditional scale recovery approaches since it does not require any fine-tuning of the monocular depth estimation neural network. Moreover, by using predictions of Depth Anything V1 whose weights are publically available, we ensure generalization capacities to a large variety of image domains. Finally, our method is robust to noisy reference points or disparity maps, allowing the use of low-quality sensors for metric depth estimation at the lowest cost. To corroborate our claims we carry out experiments on standard depth estimation benchmarks that show that our approach is competitive with other zero-shot monocular depth estimation methods. Our future works will focus on confirming the interest of our approach by comparing it with a high-resolution LiDAR in the context of off-road navigation with a real robot.

### v2

@@ IV-C Comparison with monocular depth esimation methods
Our w/ SFM (SIFT) 0.776 0.161 5.929
Our w/ SFM (Omniglue) 0.947 0.112 5.300
In Tab. I, we compare our LiDAR-based rescaling approach with zero-shot monocular depth estimation methods and other depth estimation methods that have been fine-tuned on the NYUv2 dataset [31]. We conducted a similar study in Tab. II on the KITTI dataset [35] with in addition results of rescaling with 3D reference points provided by structure-from-motion. Tab. III and Tab. IV provide zero-shot performance on indoor and outdoor datasets, respectively. For each dataset, our results show rescaling using LiDARs with different numbers of beams. Additionally, we present results of rescaling with structure-from-motion for the DDAD dataset [36]. We do not provide results of rescaling with structure-from-motion on NYUv2 [31], SUN-RGBD [32], IBIMS-1 [33] and DIODE Indoor and Outdoor [34] because they do not consist of temporal image sequences with poses. The results of the monocular depth estimation methods we compare against were directly taken from their respective papers, which explain missing numbers when they have not been reported by their authors.
Rescaling with LiDAR
Our experiments highlight an overall benefit of using our rescaling approach with $32$-laser or $16$-laser LiDAR rather than using zero-shot monocular metric depth estimation. Thus, we observe on average $46$% improvement on $\delta_{1}$, $0.5$% on AbsRel and $47$% on RMS relative to the other zero-shot methods. In contrast, the methods that have been fine-tuned on the same domain as the test set (see lines with (FT) in Tab. I and Tab. II) compare favorably to ours. These approaches benefit from an additional in-domain training which is often costly due to dataset creation and training computation. We notice that our method is sensitive to the domain with on average $28$% and $13$% enhancements on indoor (including NYUv2) and outdoor (including KITTI) datasets, respectively, relative to other zero-shot methods whatever the metric. Regarding performance with 2D LiDAR, the same sensitivity to the domain is apparent. However, the results are a bit lower than zero-shot and fine-tuned metric depth estimation methods and lower than our other approaches, especially on outdoor datasets while remaining competitive on indoor datasets. Interestingly, we observe that increasing the number of beams from 16 to 32 does not necessarily improve the performance.
Rescaling with structure-from-motion
Our experiments highlight that structure-from-motion techniques can provide reliable reference points for rescaling Depth Anything affine-invariant disparity predictions. More specifically, we show that results with SIFT [37] are slightly lower than the ones of other zero-shot monocular metric depth estimation methods, while using OmniGlue [38] increases performance relative to other zero-shot methods of the $\delta_{1}$, AbsRel and RMS metrics on average by $7$%, $4$% and $19$%, respectively. Compared to the other approaches, rescaling with SFM performs better than rescaling with LiDAR 1 beam but worse than other methods that may be more complex to set up including rescaling with LiDAR 16 and 32 beams or fine-tuned approaches. However, such approaches require LiDAR camera calibration or an additional training on a dataset that needs be created in real cases.
IV-D Comparison with depth completion methods
TABLE V: Quantitative comparison with depht completion methods on different zero-shot outdoor benchmarks.
Method DIODE Outdoor [34] DDAD [36]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
CompletionFormer [8] w/ LiDAR 64 beams 0.679 0.439 6.323 0.711 0.173 8.612
[…]
TABLE VI: Quantitative comparison with depth completion methods on different zero-shot indoor benchmarks.
Method SUN-RGBD [32] IBIMS-1 [33] DIODE Indoor [34]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
CompletionFormer [8] w/ 500 random samples 0.964 0.129 0.243 0.967 0.080 0.315 0.958 0.252 0.403
[…]
Figure 2: Quantitative study of the impact of the number of LiDAR beams on the performance of depth completion method on KITTI dataset [35].
Figure 3: Quantitative study of the impact of the number of depth samples on the performance of depth completion method on NYUv2 dataset [31].
Figure 4: Quantitative study of the impact of random noise on depth samples on the performance of depth completion method on NYUv2 dataset [31].
In this section, we compare our method with two recent depth completion approaches, CompletionFormer [8] and BP-Net [9]. Both methods have been trained once on KITTI [35] and once on NYUv2 [31]. For the sake of fairness, we conducted our evaluations with the settings used to train CompletionFormer [8] and BP-Net [9]. Thus, for NYUv2 and other indoor benchmarks, images are resized to $320\times 240$ and then center-cropped to $304\times 228$, the sparse depth is obtained by randomly sampling 500 points in the ground truth depth. For KITTI and other outdoor benchmarks, images are cropped to $1216\times 256$ and the number of LiDAR beams is set to 64. We note that no evaluation of BP-Net [9] is performed on SUN-RGBD [32], DIODE Indoor and DIODE Outdoor [34] since the camera calibration of the input images that is required by BP-Net is absent in those datasets.
Zero-shot depth completion
We evaluate zero-shot performance on outdoor and indoor benchmarks in Tab. V and Tab. VI with the depth completion networks trained on KITTI [35] and NYUv2 [31], respectively. For outdoor datasets, the results show that our method performs on par with CompletionFormer [8] on DIODE Outdoor [34] and is better on DDAD [36] for all metrics while BP-Net [9] appears to be advantageous on the latter benchmark. Regarding indoor datasets, performance on IBIMS-1 [33] and DIODE indoor are close. The fact that CompletionFormer outperforms our method on SUN-RGBD may be partially explained by very similar domains with NYUv2 [31] (used to train CompletionFormer) since one of the cameras used in SUN-RGBD is the same as that of NYUv2.
Robustness to reduction of the sparse depth density
The cases previously analyzed represent ideal conditions, as the test settings precisely match those of the training stage. However, in real cases, test conditions may be imposed or may change. If they differ significantly from those used during the training of any available depth completion model, retraining may be necessary to ensure reliable performance. We illustrate this point by examining the robustness of depth completion methods to changes in the distribution of the sparse depth. Indeed, in practical scenarios, the sparse depth distribution is given by the type of LiDAR used. Additionally, when SFM is employed, variations in image texture can cause fluctuations in sparse depth density over time. Fig. 2 and Fig. 3 show the influence of decreasing the number of LiDAR beams (from 64 to 1) and the number of points in the sparse depth (from 500 to 10) on the KITTI Depth Completion [39] and NYUv2 [31] datasets, respectively. The results highlight the robustness of our method to the reduction of the sparse depth density relative to depth completion techniques, as their performance drops when ours is barely affected. The advantage of our method comes from its simplicity since only two 3D points are necessary in theory to regress the two rescaling parameters. However, depth completion robustness is likely to be improved by randomly varying the density of the sparse depth at training. We note that the better results of depth completion methods on a large number of samples are likely to be due to the evaluation sets which come from the same datasets as their training data contrary to our approach.
Robustness to noisy sparse depth
Another limitation of the analysis we carry out on zero-shot depth completion is that the sparse depth corresponds to the ground truth while in real cases, it may be noisy. To assess the robustness of depth completion methods to noise, we conduct experiments on the NYUv2 dataset [31] where a centered Gaussian noise is added to the inputs sparse depth. We study the impact of gradually increasing its standard deviation from 5cm to 1 meter. We have chosen Gaussian noise for its simplicity but other types of noise commonly encountered in SFM or LiDAR could have been considered. The results in Fig. 4 demonstrate the robustness of our approach relative to noisy depth samples. Unlike other depth completion approaches, which suffer from significant performance degradation as noise increases, our method remains stable. However, the robustness to noisy sparse depth could be improved by adding random noise at training stage as data augmentation. Again, we note that the better results of depth completion methods with low-noise level is likely to be due to overfitting on the NYUv2 dataset.
IV-E Inference cost study
TABLE VII: Cost comparison for different architectures.
(ZS), (FT) and (DC) means Zero-Shot, Fine-Tuned and Depth Completion, respectively.
@@ IV-E Inference cost study
CompletionFormer [8] (DC) 84 150
BP-Net [9] (DC) 90 103
Ours (Depth Anything [2] + rescaling) 335 120
@@
To complete our study, we analyze the inference cost of rescaling Depth Anything V1 affine-invariant disparity maps with the other zero-shot (ZS), fine-tuned (FT) monocular depth estimation or depth completion (DC) methods in Tab. VII. For this purpose, we compare two informative and easy-to-access variables which are the number of parameters and the average inference runtime of these approaches for a single image. Note that the runtimes have been measured on the same NVIDIA GeForce RTX $3090$ GPU for each approach. We notice that zero-shot methods are more costly at inference than our rescaling approach because they introduce extra modules such as a ConvGRU block in [30] or a camera module in [7] that returns a dense representation of the camera calibration. In contrast, fine-tuned methods have lower inference costs relative to ours thanks to their more complex training. As for depth completion methods, they are always lighter in terms of parameters which can be justified for architectures designed to be trained on a single dataset. Consequently, we could have expected a low inference time, which is what we observe with BP-Net but not with CompletionFormer [8]. It may be due to exotic modules in their architecture such as the Joint Convolutional Attention and Transformer block (JCAT). However, one may think that CompletionFormer could be sped up by running independent modules in parallel. As each image is processed independently, the computation overhead of our method can be even reduced when applied to a video stream. Indeed, since the neural network runs on the GPU and the RANSAC on the CPU, the inference of the neural network at step $t+1$ can overlap the execution of RANSAC at step $t$.
IV-F Comparison with Depth Anything V2
@@
TABLE VIII: Comparative study between Depth Anything V1 and Depth Anything V2 on the KITTI dataset [35].
Rescaling with Depth Anything KITTI [35]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$ $R^{2}$uparrow
@@
We compare performance between Depth Anything V1 [2] and Depth Anything V2 [5] on the KITTI dataset in Tab. VIII. We notice that Depth Anything V1 [2] slightly outperforms Depth Anything V2 [5] when used in our rescaling approach. Moreover, Depth Anything V1 has a higher coefficient of determination $R^{2}$ which corresponds to the proportion of the variance of the metric disparity that is explainable by the linear regression model parameterized with the $\alpha$ and $\beta$ that has been found. This means that while Depth Anything V2 manages to handle small details better than Depth Anything V1 (see Fig. 5), the latter predicts more accurate disparity maps within an affine transformation.
@@
IV-G Ablation study
TABLE IX: Ablation study to validate the need for estimating the rescaling parameters for each image
Rescaling with NYUv2 [31]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
@@ IV-G Ablation study
KITTI [35]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
@@
[…]
TABLE X: Ablation study to validate the use of the RANSAC.
Methods RANSAC KITTI [35]
$\delta_{1}$uparrow AbsRel $\downarrow$ RMS $\downarrow$
Ours w/ LiDAR 1 beam 0.897 0.113 4.871
@@
Ours w/ LiDAR 16 beams 0.966 0.063 2.849
@@
Ours w/ LiDAR 32 beams 0.966 0.063 2.852
@@
Tab. IX compares dynamic rescaling, i.e., an estimation of the parameters $\alpha$ and $\beta$ for each image to a static rescaling with unique parameters for all the images of a dataset. In the latter case, the parameters are the mean $\alpha$ and $\beta$ obtained with the best rescaling (here with LiDAR 16 beams). The results show that performing a rescaling for each image tends to provide better performance even if the camera does not change or if the image domain remains similar.
The ablation study in Tab. X aims to justify the use of a RANSAC algorithm [11] when estimating the parameters of the affine transformation to recover the metric depth. We observe that the RANSAC is beneficial most of the time, especially when the 3D reference points are not provided by a LiDAR. Indeed, structure-from-motion are more likely to generate outliers that would disturb a vanilla linear regression but would be filtered out by the RANSAC.
@@ V Conclusion
In this paper, we provide a straightforward method to estimate monocular metric depth by rescaling Depth Anything V1 [2] affine-invariant disparity predictions using 3D reference points provided by an external sensor or technique. By using Depth Anything V1 predictions whose weights are publicly available, we ensure generalization capacities to a large variety of image domains. By leveraging a RANSAC, our method is robust to noise in the sparse depth or in the disparity maps, allowing the use of low-quality sensors for metric depth estimation. Thus, the solution we propose is adaptable to any camera calibration and does not need any fine-tuning of the monocular depth estimation neural network, which in practice also means that no costly creation of a dataset of the target domain is required. For all these reasons, our approach may also be a good candidate for providing monocular metric depth at low cost. To corroborate our claims we carry out experiments on standard depth estimation benchmarks that show that our approach is competitive with other zero-shot monocular depth estimation methods. We also demonstrate superiority with respect to depth completion methods in downgraded mode. Our future works will focus on confirming the advantages of our approach by comparing it with a high-resolution LiDAR in the context of off-road navigation with a real robot.